OpenStudy (anonymous):

Soren has two jobs. During the day he works as an office clerk. In the evening he works as a cashier. His office job pays him $11 per hour. His cashier job pays him $8.75 per hour. In one week, Soren worked 46 hours. He earned a total of $470. Let c represent the number of hours that Soren works as an office clerk. Which equation can be used to find how many hours Soren worked in both jobs?

OpenStudy (anonymous):

A. 11c + 8.75(46 – c) = 470 B. 11c + 8.75(c – 46) = 470 C. 11c + 8.75(470 – c) = 46 D. 8.75c + 11(46 – c) = 470